# TacEleven: generative tactic discovery for football open play

### v1

@@ 2 Result
To evaluate the effectiveness and explainability of TacEleven, we design three tasks with progressive tactical complexity: counterfactual exploration (CF), single-step discovery ($S$), and multi-step discovery (MS). Each task combines both qualitative and quantitative analyses to ensure reliable and interpretable evaluation.
@@ 3 Refer to caption Refer to caption Refer to caption
Quantitative analysis. We quantitatively evaluate the improvement of the discovered single-step tactics in the $S$ task of Table 1. According to player roles, we separately report the results for players in the front line (Front), midfield (Middle), back line (Back), and for the entire team (Entire). Overall, the single-step tactics discovered by TacEleven outperform the Factual in all metrics except consistency, indicating that the model successfully discovers superior single-step tactics. Compared with the Random, the consistency improves by 8%, showing that MTC not only identifies better tactics in a transparent and explainable manner but also filters out unreasonable ones. Across different field zones, the performance metrics exhibit distinct patterns. The Front shows the most significant improvement, suggesting that in the attacking third, TacEleven can more effectively mine tactics that create offensive advantages.
@@ 4.1 Problem Formulation
Tactic definition. To model tactics in open play, we first define a set of meta-actions $\mathbbm{A}$, the most fundamental and atomic actions in football. These meta-actions are both minimal and complete, capturing the essence of individual events (e.g., pass, carry, shot, etc.) while forming the building blocks for more complex tactical patterns. By combining different meta-actions $a_{1},a_{2},\ldots\in\mathbbm{A}$, a tactic $\mathbf{T}$ can be formed as $\langle a_{1},a_{2},\ldots,a_{N}\rangle$, where NN denotes the step number of the meta-actions, introducing abundant fine-grained features. Compared to relying solely on semantic descriptions, the incorporation of these fine-grained features offers more precise and detailed low-level representations, enabling a deeper understanding of tactical structures.
Spatiotemporal graph representing. The spatiotemporal graph (STG) offers a powerful framework for modeling these relationships, combining spatial and temporal features with semantic descriptions. This representation not only facilitates the integration of multi-modal data but also enhances the model’s ability to capture the complexities of football tactics, making it a cornerstone of our approach. A STG $\mathcal{G}$ can comprehensively represent $\mathbbm{A}$. The STG $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X},A,T,s)$ is assumed to be static, with nodes $\mathcal{V}$, edges $\mathcal{E}\in\mathcal{V}\times\mathcal{V}$, node feature sequence $\mathbf{X}$, a stable adjacency matrix AA, the temporal index $T$ of $\mathbf{X}$, and an global language description $s$. Within the context of football games, $\mathcal{V}$ is defined as the set of 23 entities (22 players and 1 ball) present on the pitch, and $\mathcal{E}$ is constructed as the set of all possible pairwise connections among them. Furthermore, the STG is assumed to encode relevant features. In particular, a node feature sequence $\mathbf{X}\in\mathbbm{R}^{|T|\times|\mathcal{V}|\times d}$ with feature dimension $d$, a adjacency matrix $A\in\mathbbm{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ with binary elements, and a language description $s$ are provided, where $|T|$ denotes the length of the temporal sequence, and $|V|$ denotes the number of nodes. The relevant entries of these objects serve as input features for each node and edge. The $\mathbf{X}_{v}\in\mathbbm{R}^{|T|\times d}$ represents the attributes of an individual player or the ball $v\in\mathcal{V}$, (e.g., position, velocity, and role). The edge $A_{uv}\in\{0,1\}$ represents the attribute of a particular pair of entities $(u,v)\in\mathcal{E}$, (e.g., whether two players are teammates). The language description $s$ is a text segment that characterizes the corresponding tactic instance (e.g., event type, participating players, etc.).
@@ 4.2 Data Preprocessing
Furthermore, to construct the prepared $\mathcal{G}$, the raw data are filtered after the timestamp alignment, due to incomplete information or identifier inconsistencies. After filtering, 1,076,258 valid meta-actions remained for training and evaluation. Subsequently, the node feature sequence $\mathbf{X}$ is extracted from the tracking trajectory data within $|T|=5$ by uniform temporal sampling. the adjacency matrix AA encodes as an all-ones matrix representing a fully connected graph. Finally, rich tactical description annotations, integrated from the event data, are appended to these instances.
@@ 4.3 Language-controlled Tactical Generator Details
Spatiotemporal attention. In our model, the attention mechanism [attention] is incorporated to effectively capture spatiotemporal dependencies and align multi-modal inputs, thereby strengthening the model’s predictive performance. An attention mechanism is determined by a similarity function between the query matrix $Q$ and the key matrix $K$, and the resulting weighted combination is applied to the value matrix $V$. In practice, these matrices are projected into designated dimensions with multiple axes, typically structured as $Q,K,V\in\mathbbm{R}^{N\times|T|\times|\mathcal{V}|\times D}$, where NN is the batch size. We define the spatiotemporal attention with as $\mathrm{Attention}^{\mathrm{ST}}(X^{Q},X^{K},X^{V})$, with three input matrices $X^{Q},X^{K},X^{V}\in\mathbbm{R}^{N\times|T|\times|\mathcal{V}|\times d_{X^{*}}}$, where $d_{X^{*}}$ respectively denotes the feature dimension of them. The non-linear projections separately transform the input matrices:
@@ 4.4 Tactical Tree Search
The tactical tree search is motivated by the need for a systematic approach to tactical discovery in football games. Tree search methods are widely adopted in AI-powered scientific discovery for efficiently exploring large combinatorial spaces [segler2018planning, Dieb_Ju_Shiomi_Tsuda_2019, wang2023scientific, yamada2025aiscientistv2workshoplevelautomated]. By modeling the process as a tree search, we can effectively explore the counterfactual space and identify optimal tactical sequences that align with high-level instructions. The MTC leverages an MLLM $\mathcal{M}$ (QVQ-Max in practice) to accomplish the tactic discovery task. Whether for single-step or multi-step tactical discovery, we uniformly model the process as a tactical tree search. As illustrated in Fig. 8, the root node represents the historical context $\mathcal{G}_{0}$, the degree of each node is the size of the counterfactual space $|\mathcal{D}_{n}|$, and the depth of the tree corresponds to the predefined steps NN. In the process of tactic discovery, $\mathcal{M}$ starts from the root node and searches for a path such that the formed meta-action sequence best aligns with the instruction $P$. Leveraging the planning and reasoning capabilities of MLLMs, $\mathcal{M}$ as $f_{\psi}(\cdot)$ is capable of making optimal search decisions at each step. We meticulously designed prompts to encompass the entire tactic discovery pipeline along with a tactical style pool, as represented in the Appendix E.

### v2

@@ 2 Result
To evaluate the effectiveness and explainability of TacEleven, we design three tasks with progressive tactical complexity: counterfactual exploration (CF), single-step discovery (SS), and multi-step discovery (MS). Each task combines both qualitative and quantitative analyses to ensure reliable and interpretable evaluation.
@@ 3 Refer to caption Refer to caption Refer to caption
Quantitative analysis. We quantitatively evaluate the improvement of the discovered single-step tactics in the SS task of Table 1. According to player roles, we separately report the results for players in the front line (Front), midfield (Middle), back line (Back), and for the entire team (Entire). Overall, the single-step tactics discovered by TacEleven outperform the Factual in all metrics except consistency, indicating that the model successfully discovers superior single-step tactics. Compared with the Random, the consistency improves by 8%, showing that MTC not only identifies better tactics in a transparent and explainable manner but also filters out unreasonable ones. Across different field zones, the performance metrics exhibit distinct patterns. The Front shows the most significant improvement, suggesting that in the attacking third, TacEleven can more effectively mine tactics that create offensive advantages.
@@ 4.1 Problem Formulation
Tactic definition. To model tactics in open play, we first define a set of meta-actions $\mathbbm{A}$, the most fundamental and atomic actions in football. These meta-actions are both minimal and complete, capturing the essence of individual events (e.g., pass, carry, shot, etc.) while forming the building blocks for more complex tactical patterns. By combining different meta-actions $a_{1},a_{2},\ldots\in\mathbbm{A}$, a tactic $\mathbf{T}$ can be formed as $\langle a_{1},a_{2},\ldots,a_{N}\rangle$, where $N$ denotes the step number of the meta-actions, introducing abundant fine-grained features. Compared to relying solely on semantic descriptions, the incorporation of these fine-grained features offers more precise and detailed low-level representations, enabling a deeper understanding of tactical structures.
Spatiotemporal graph representing. The spatiotemporal graph (STG) offers a powerful framework for modeling these relationships, combining spatial and temporal features with semantic descriptions. This representation not only facilitates the integration of multi-modal data but also enhances the model’s ability to capture the complexities of football tactics, making it a cornerstone of our approach. A STG $\mathcal{G}$ can comprehensively represent $\mathbbm{A}$. The STG $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X},A,T,s)$ is assumed to be static, with nodes $\mathcal{V}$, edges $\mathcal{E}\in\mathcal{V}\times\mathcal{V}$, node feature sequence $\mathbf{X}$, a stable adjacency matrix $A$, the temporal index $T$ of $\mathbf{X}$, and an global language description $s$. Within the context of football games, $\mathcal{V}$ is defined as the set of 23 entities (22 players and 1 ball) present on the pitch, and $\mathcal{E}$ is constructed as the set of all possible pairwise connections among them. Furthermore, the STG is assumed to encode relevant features. In particular, a node feature sequence $\mathbf{X}\in\mathbbm{R}^{|T|\times|\mathcal{V}|\times d}$ with feature dimension $d$, a adjacency matrix $A\in\mathbbm{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ with binary elements, and a language description $s$ are provided, where $|T|$ denotes the length of the temporal sequence, and $|V|$ denotes the number of nodes. The relevant entries of these objects serve as input features for each node and edge. The $\mathbf{X}_{v}\in\mathbbm{R}^{|T|\times d}$ represents the attributes of an individual player or the ball $v\in\mathcal{V}$, (e.g., position, velocity, and role). The edge $A_{uv}\in\{0,1\}$ represents the attribute of a particular pair of entities $(u,v)\in\mathcal{E}$, (e.g., whether two players are teammates). The language description $s$ is a text segment that characterizes the corresponding tactic instance (e.g., event type, participating players, etc.).
@@ 4.2 Data Preprocessing
Furthermore, to construct the prepared $\mathcal{G}$, the raw data are filtered after the timestamp alignment, due to incomplete information or identifier inconsistencies. After filtering, 1,076,258 valid meta-actions remained for training and evaluation. Subsequently, the node feature sequence $\mathbf{X}$ is extracted from the tracking trajectory data within $|T|=5$ by uniform temporal sampling. the adjacency matrix $A$ encodes as an all-ones matrix representing a fully connected graph. Finally, rich tactical description annotations, integrated from the event data, are appended to these instances.
@@ 4.3 Language-controlled Tactical Generator Details
Spatiotemporal attention. In our model, the attention mechanism [attention] is incorporated to effectively capture spatiotemporal dependencies and align multi-modal inputs, thereby strengthening the model’s predictive performance. An attention mechanism is determined by a similarity function between the query matrix $Q$ and the key matrix $K$, and the resulting weighted combination is applied to the value matrix $V$. In practice, these matrices are projected into designated dimensions with multiple axes, typically structured as $Q,K,V\in\mathbbm{R}^{N\times|T|\times|\mathcal{V}|\times D}$, where $N$ is the batch size. We define the spatiotemporal attention with as $\mathrm{Attention}^{\mathrm{ST}}(X^{Q},X^{K},X^{V})$, with three input matrices $X^{Q},X^{K},X^{V}\in\mathbbm{R}^{N\times|T|\times|\mathcal{V}|\times d_{X^{*}}}$, where $d_{X^{*}}$ respectively denotes the feature dimension of them. The non-linear projections separately transform the input matrices:
@@ 4.4 Tactical Tree Search
The tactical tree search is motivated by the need for a systematic approach to tactical discovery in football games. Tree search methods are widely adopted in AI-powered scientific discovery for efficiently exploring large combinatorial spaces [segler2018planning, Dieb_Ju_Shiomi_Tsuda_2019, wang2023scientific, yamada2025aiscientistv2workshoplevelautomated]. By modeling the process as a tree search, we can effectively explore the counterfactual space and identify optimal tactical sequences that align with high-level instructions. The MTC leverages an MLLM $\mathcal{M}$ (QVQ-Max in practice) to accomplish the tactic discovery task. Whether for single-step or multi-step tactical discovery, we uniformly model the process as a tactical tree search. As illustrated in Fig. 8, the root node represents the historical context $\mathcal{G}_{0}$, the degree of each node is the size of the counterfactual space $|\mathcal{D}_{n}|$, and the depth of the tree corresponds to the predefined steps $N$. In the process of tactic discovery, $\mathcal{M}$ starts from the root node and searches for a path such that the formed meta-action sequence best aligns with the instruction $P$. Leveraging the planning and reasoning capabilities of MLLMs, $\mathcal{M}$ as $f_{\psi}(\cdot)$ is capable of making optimal search decisions at each step. We meticulously designed prompts to encompass the entire tactic discovery pipeline along with a tactical style pool, as represented in the Appendix E.